# A Copula-Based Variational Autoencoder for Uncertainty Quantification in Inverse Problems: Application to Damage Identification in an Offshore Wind Turbine

### v1

@@ 2.2 Synthetic data generation
In this work, we employ the NREL’s open-source wind turbine simulation tool OpenFAST ([43]), which accounts for the influence of aerodynamic ([52, 53]) and hydrodynamic ([54]) excitations on the response of the floating platform. OpenFAST is a highly accurate and reliable tool to perform numerical simulations of FOWTs [55, 56, 57].
@@
2.3 Loss function for the inverse problem
When damage occurs affecting the mooring system of the FOWT, its motion response will change with respect to that observed under healthy conditions. Identifying the mooring system’s state based on measurements of its motion response (and environmental excitation) is an inverse problem, $\mathcal{I}$, which maps the observations into a set of damage descriptors or features, $\mathbf{z}\in\mathcal{Z}$, such that: $\mathcal{I}:\mathcal{M}\times\mathcal{W}\rightarrow\mathcal{Z}$, where $\mathbf{z}$ contains $D\geq 2$ features.
Let us assume that the measured motion responses, environmental conditions (wind and wave excitation), and the unknown system condition are random variables. Vectors $\mathbf{m}$, $\mathbf{w}$, and $\mathbf{z}$ now correspond to realizations of each of these sets of random variables. Considering the existence of additive noise that is inherent to the data acquisition process, we express the forward problem as:
@@ 2.3 Loss function for the inverse problem
where the damage condition $\mathbf{z}$ and the environmental (wind and wave) excitation $\mathbf{w}$ prescribe the response of the system according to the exact forward operator $\mathcal{F}$ plus an additive noise term $\bm{\epsilon}$
In [29], we designed a Bayesian VAE architecture to statistically describe the posterior distribution of the estimated damaged properties $\mathbf{z}$, given measurements of the system’s response $\mathbf{m}$ and environmental conditions $\mathbf{w}$. The architecture includes an encoder $\mathcal{I}_{\bm{\theta}}$ to map the response measurements and environmental conditions to the parameters ($\mathbf{p}$) of a parametric posterior Probability Density Function (PDF) denoted as $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$. Next, it embeds a sampling layer —adapted to the selected posterior PDF — to draw random samples that enter the decoder, $\mathcal{F}_{\bm{\varphi}}$. The decoder approximates the exact forward operator $\mathcal{F}$ via a fully-connected NN with parameters $\bm{\varphi}$. Figure 3 describes the proposed architecture.
@@
Figure 3: Variational Autoencoder architecture. The encoder estimates the properties that describe the posterior PDF of the latent space (damaged condition features), $\mathbf{z}\sim q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$. A sampling layer draws $H$ random samples from the distribution, which are then fed to the optimal forward operator $\mathcal{F}_{\bm{\varphi}^{*}}$. The output layer yields the reconstruction of the input measurements, $\hat{\mathbf{m}}_{i}^{h}=\mathcal{F}_{\bm{\varphi}^{*}}\circ\mathcal{I}_{\bm{\theta}}(\mathbf{m}_{i}^{h},\mathbf{w}_{i})$, for each sample h=1,,Hh=1,...,H.
The training phase requires two steps [23, 28, 14, 29]. The decoder $\mathcal{F}_{\bm{\varphi}}$ is first trained to obtain the optimal parameters $\bm{\varphi}^{*}$. See previous works [14, 29] for details on this deterministic training step. Next, keeping the optimal decoder parameters ($\bm{\varphi}^{*}$), we train the entire architecture to obtain $\bm{\theta}^{*}$. After completing the training, for any measured input sample $\mathbf{m}$, we obtain the parameters describing the approximate posterior PDF of the damage condition properties, $\mathbf{z}$. The contour plot of the estimated PDFs reveals the regions in the damage condition space $\mathcal{Z}$ that are more likely to have produced the observed measurements $\mathbf{m}$.
In order to define the loss function required for the second training step (learning the inverse operator $\mathcal{I}_{\bm{\theta}}$), we first substitute $\mathcal{F}$ in Eq. 11 by its optimal approximation $\mathcal{F}_{\bm{\varphi}^{*}}$. We assume that the unknown measurement error $\bm{\epsilon}$ follows a known PDF ($\bm{\epsilon}\sim p(\bm{\epsilon})$) and that it is mutually independent with respect to the unknown damage condition $\mathbf{z}$ [26]. Thus, according to Eq. 11, we can express the likelihood as:
$p(\mathbf{m}|\mathbf{w},\mathbf{z})=p\left(\mathbf{m}-\mathcal{F}_{\bm{\varphi}^{*}}([\mathbf{z},\mathbf{w}])\right).$ (12)
In this work, we assume the noise follows a Gaussian distribution, $p(\bm{\epsilon})=\mathcal{N}(0,\Gamma)$, where $\Gamma=\text{diag}(\beta\mathcal{F}_{\bm{\varphi}^{*}}([\mathbf{z},\mathbf{w}]))^{2}$ is a vector that contains the non-zero elements of a diagonal matrix representing the variances of the noise, and the parameter $\beta$ corresponds to the noise level. This allows us to rewrite Eq. 12 as:
@@
Although $\mathbf{z}$ is unknown, we can represent its uncertainty using a conditional probability distribution considering its relationship with the measured variables, $p(\mathbf{z}|\mathbf{m},\mathbf{w})$. For any damage condition $\mathbf{z}$, the conditional PDF $p(\mathbf{z}|\mathbf{m},\mathbf{w})$ is the target posterior distribution that solves the inverse problem $\mathcal{I}$ under the Bayesian framework. We employ Bayes’ theorem [58] to express it by the following proportionality:
@@
Due to the intractability of the true posterior (usually an unknown non-parametric PDF), we define an inverse operator $\mathcal{I}_{\bm{\theta}}$ that estimates the parameters of an approximate PDF for each damage condition in $\mathbf{z}$, given the motion measurements $\mathbf{m}$ and operating loadings $\mathbf{w}$.
Compared to the deterministic approach, instead of a single value estimate, now the output of the inverse operator $\mathcal{I}_{\bm{\theta}}$ spans to produce the set of properties describing the posterior PDF. Realizations from the approximate posterior $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ represent samples of the damage condition $\mathbf{z}$ that are likely produced by the unknown true posterior $p(\mathbf{z}|\mathbf{m},\mathbf{w})$. All the realizations $\{\mathbf{z}_{h}\}_{h=1}^{H}$ share the same operating conditions $\mathbf{w}$.
In this variational approach, we aim at maximizing the likelihood of the data while at the same time minimizing the discrepancy between the true posterior $p(\mathbf{z}|\mathbf{m},\mathbf{w})$ and the approximate posterior $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ obtained from $\mathcal{I}_{\bm{\theta}}$ The commonly employed loss function to achieve this goal is the Evidence Lower BOund (ELBO) [59]. For any distributional model $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$, we can express the evidence as [24]:
$\begin{split}\log p(\mathbf{m},\mathbf{w})=\mathbb{E}_{q_{\bm{\theta}}}[\log p(\mathbf{m},\mathbf{w})]&=\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{p(\mathbf{m},\mathbf{w}|\mathbf{z})}{p(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]\\ &=\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{p(\mathbf{m},\mathbf{w}|\mathbf{z})q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})p(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]\\ &=\underbrace{\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{p(\mathbf{m},\mathbf{w}|\mathbf{z})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{ELBO}}}+\underbrace{\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})}{p(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}D_{\text{KL}}(q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})||p(\mathbf{z}|\mathbf{m},\mathbf{w})}},\end{split}$ (15)
where the Kullback-Leibler divergence $D_{\text{KL}}$ is a statistical measure of the distance between two PDFs. It achieves a null value when the compared PDFs are equal and can be defined as [24]:
$D_{\text{KL}}[p(x)||q(x)]=\int p(x)\ \text{log}\frac{p(x)}{q(x)}dx,$ (16)
where $x$ indicates a realization of the random variable $X$, and $p(x)$ and $q(x)$ are the two different PDFs. The use of logarithms aids in avoiding numerical instabilities [60]. Since $D_{\text{KL}}$ is non-negative by definition, the first term in 15 constitutes a lower bound. Rearranging terms, we define the ELBO loss as:
@@
where we assume that the operating conditions $\mathbf{w}$ and the damage properties $\mathbf{z}$ are independent (i.e., $p(\mathbf{w},\mathbf{z})=p(\mathbf{w})\cdot p(\mathbf{z})$). Hence, we can get rid of $p(\mathbf{w})$ in the second line of Eq. 17 as it is independent of $\mathbf{z}$. For a certain observation $\{\mathbf{m},\mathbf{w}\}$, we draw $H$ samples from the posterior and approximate the ELBO loss function as
@@
By substituting the likelihood from Eq. 13 into Eq. 18, we finally express $\mathcal{L}_{\text{ELBO}}$ as:
@@
Here, the first term accounts for the data misfit, which is the error between the true measurements and the reconstructions provided by $\mathcal{F}_{\bm{\varphi}^{*}}$. The second term refers to the prior, which we assume to follow a bounded uniform distribution $p(\mathbf{z})\sim\mathcal{U}[\mathbf{b}_{low},\mathbf{b}_{up}]$ with lower and upper bounds $\mathbf{b}_{low}$ and $\mathbf{b}_{up}$, respectively. The last term measures the probability that the $h$-th sample belongs to the estimated distribution $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$. The second term can be neglected by directly constraining the Gaussian mixture density function to the desired interval. Note that maximizing the variational lower bound is equivalent to minimizing its negative.
@@
The next subsection explores different configurations for the approximate posterior $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ involved in Eq.20, including the necessary sampling strategy (see Figure 3).
@@ 2.4.1 Gaussian Mixture
Following our previous work [61], we start considering a Gaussian Mixture as the approximate posterior $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$, such that:
@@
$\begin{gathered}q^{\text{GM}}_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w}):=\sum_{k=1}^{K}\alpha_{k}\mathcal{N}^{\tau}(\mathbf{z}|\bm{\mu}_{k},\Sigma_{k},\Omega_{k});\\ \mathcal{N}^{\tau}(\mathbf{z}|\bm{\mu}_{k},\Sigma_{k},\Omega_{k})=\frac{1}{C_{k}}\mathcal{N}(\mathbf{z}|\bm{\mu}_{k},\Sigma_{k})\mathbb{I}(\mathbf{z}\in\Omega_{k}),\end{gathered}$ (21)
where the PDF $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ is a multivariate truncated Gaussian Mixture (GM) dependent on the parameters $\bm{\theta}$ of the DNN $\mathcal{I}_{\bm{\theta}}$, $\alpha_{k}$ is the weight of the $k$-th Gaussian component, and $\mathcal{N}(\mathbf{z}|\bm{\mu}_{k},\bm{\Sigma}_{k})$ is the corresponding multivariate Gaussian distribution with mean vector $\bm{\mu}_{k}$ and covariance matrix $\bm{\Sigma}_{k}=\text{diag}(\bm{\sigma}_{k})$ for k=1,,Kk=1,...,K Gaussians. The term $\mathcal{N}^{\tau}$ denotes the truncated Gaussian components, where $\Omega_{k}$ is the support or truncation region for the $k$-th component, namely $\mathbf{b}_{\text{low}},\mathbf{b}_{\text{up}}]$, $C_{k}$ is a normalizing constant that ensures each component integrates to 1 over its support $\Omega_{k}$, and $\mathbb{I}(\mathbf{z}\in\Omega_{k})$ denotes the indicator function, which is one if $\mathbf{z}$ is in the region $\Omega_{k}$ and zero otherwise. We obtain $C_{k}$ as an approximate integral value of the corresponding non-truncated Gaussian PDF using $N^{\text{MC}}$ Monte Carlo (MC) points over the region $\Omega_{k}$:
$C_{k}=\int_{\Omega_{k}}\mathcal{N}(\mathbf{z}|\bm{\mu}_{k},\Sigma_{k})d\mathbf{z}\approx\sum_{j=1}^{N^{\text{MC}}}\frac{1}{(2\pi)^{D/2}|\Sigma_{k}|^{1/2}}\exp{\left(-\frac{1}{2}(\mathbf{z}_{j}-\bm{\mu}_{k})^{T}\Sigma_{k}^{-1}(\mathbf{z}_{j}-\bm{\mu}_{k})\right)}d\mathbf{z}_{j}$
where the PDF $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ is a multivariate Gaussian Mixture (GM) dependent on the parameters $\bm{\theta}$ of the DNN $\mathcal{I}_{\bm{\theta}}$, $\alpha_{k}$ is the weight of the $k$-th Gaussian component, and $\mathcal{N}(\mathbf{z}|\bm{\mu}_{k},\bm{\Sigma}_{k})$ is the corresponding multivariate Gaussian distribution with mean vector $\bm{\mu}_{k}$ and covariance matrix $\bm{\Sigma}_{k}=\text{diag}(\bm{\sigma}_{k})$ for k=1,,Kk=1,...,K Gaussians.
@@
In our original work, we assumed diagonal covariance matrices. For any component in the mixture described in Eq.21, we formulated the covariance matrix $\bm{\Sigma}_{k}$ as $\bm{\Sigma}_{k}^{\text{diag}}=\text{diag}(\sigma_{1}^{2},...,\sigma_{Z}^{2})_{k}$, where $\sigma_{i}^{2}$ indicates the variance of the $i$th variable, and the subindex $k$ indicates the Gaussian component in the mixture (k=1,2,,Kk=1,2,...,K). For simplicity in notation, from now on we will omit the $k$ subindex to describe the covariance matrix specifications. The diagonal covariance matrix assumes a null interrelation between the involved variables $\mathbf{z}$, which enormously constrains the final shape of the mixture, hindering a fine approximation of the true solution [61].

### v2

@@ 2.2 Synthetic data generation
say here that the simulator solves the forward problem as $\mathcal{F}$ with some additional noise or discrepancy with respect to the exact operator but we assume it approximates well the exact operator. You can mention $\mathcal{F}$ as the exact and $\mathcal{F}^{\prime}$ as the simulator that produces the responses to be used. We use these data first to pre-train the decoder before moving to the training of the inverse. In this work, we employ the NREL’s open-source wind turbine simulation tool OpenFAST ([64]), which accounts for the influence of aerodynamic ([78, 79]) and hydrodynamic ([80]) excitations on the response of the floating platform. OpenFAST is a highly accurate and reliable tool to perform numerical simulations of FOWTs [67, 68, 69].
@@
2.3 Inverse problem for the FOWT mooring system condition
2.3.1 VAE architecture
When damage occurs affecting the mooring system of the FOWT, its motion response will change with respect to that observed under healthy conditions. Identifying the mooring system’s state based on measurements of its motion response $\mathbf{m}\in\mathcal{M}$ (and environmental excitation $\mathbf{w}\in\mathcal{W}$ ) is an inverse problem, $\mathcal{I}$, which maps the observations into a set of damage descriptors or features, $\mathbf{z}\in\mathcal{Z}$, such that: $\mathcal{I}:\mathcal{M}\times\mathcal{W}\rightarrow\mathcal{Z}$, where $\mathbf{z}$ contains $D\geq 2$ features.
To describe the uncertainty in the measured motion responses $\mathbf{m}$, environmental conditions (wind and wave excitation) $\mathbf{w}$, and unknown system damage condition $\mathbf{z}$, we model them as random variables. These quantities are related through the equation:
@@ 2.3.1 VAE architecture
where $\bm{\epsilon}$ is an additive noise term that represents data acquisition and modeling errors, and $\mathcal{F}$ denotes the exact forward operator that produces the motion response of the system $\mathbf{m}$.
Following [52], we design a Bayesian VAE architecture to estimate the posterior distribution of the estimated damage properties $\mathbf{z}$ given measurements of the system’s response $\mathbf{m}$ and of the environmental conditions $\mathbf{w}$. This architecture is formed by the following components (see Figure 3 for a schematic): an encoder, a sampling layer, and a decoder. The encoder is a fully-connected neural network (NN) $\mathcal{E}_{\bm{\theta}}$ parametrized by $\bm{\theta}$ that maps the motion $\mathbf{m}$ and environmental condition $\mathbf{w}$ measurements to the properties $\bm{\zeta}_{\bm{\theta}}\in\mathbb{R}^{n_{p}}$ of a parametric posterior probability density function (PDF) denoted as $q_{\bm{\zeta}_{\bm{\theta}}}(\mathbf{z}|\mathbf{m},\mathbf{w})$. Symbol $q(\cdot)$ represents a computationally tractable approximation of a true PDF, $p(\cdot)$. Particularly, $q_{\bm{\zeta}_{\bm{\theta}}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ approximates the true and unknown posterior PDF $p(\mathbf{z}|\mathbf{m},\mathbf{w})$. From now on, we express the dependence of the posterior on the estimated properties $\bm{\zeta}_{\bm{\theta}}$ directly through the NN parameters $\bm{\theta}$, i.e., $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$. The sampling layer —adapted to the selected posterior PDF — draws random samples from the parametric posterior, which correspond to the damage condition properties. The decoder $\mathcal{F}_{\bm{\varphi}^{*}}$ is a pretrained fully-connected NN parametrized by $\bm{\varphi}^{*}$ that reconstructs the system’s motion response. It acts as a surrogate that substitutes for the exact forward operator $\mathcal{F}$ or its computationally expensive OpenFAST solver $\mathcal{F}^{\prime}$, enabling the timely running of simulations required to learn the optimal encoder. Details on the training step to obtain the optimal decoder NN ($\mathcal{F}_{\bm{\varphi}^{*}}$) can be found in [81, 52].
@@
Figure 3: correct figure notation for the encoder outputs Variational Autoencoder architecture. The encoder estimates the properties that describe the posterior PDF of the latent space (damaged condition features), $\mathbf{z}\sim q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$. A sampling layer draws $H$ random samples from the distribution, which are then fed to the optimal forward operator $\mathcal{F}_{\bm{\varphi}^{*}}$. The output layer yields the reconstruction of the input measurements, $\hat{\mathbf{m}}_{i}^{h}=\mathcal{F}_{\bm{\varphi}^{*}}\circ\mathcal{E}_{\bm{\theta}}(\mathbf{m}_{i}^{h},\mathbf{w}_{i})$, for each sample h=1,,Hh=1,...,H.
2.3.2 Loss function
To train the encoder NN ($\mathcal{E}_{\bm{\theta}}$), we freeze the optimal decoder parameters ($\bm{\varphi}^{*}$) and train the entire VAE to learn the optimal parameter set $\bm{\theta}=\bm{\theta}^{*}$. After completing the training, for any measured input sample $\mathbf{m}$, we obtain the set of properties $\bm{\zeta}_{\bm{\theta}^{*}}$ parametrizing the approximate posterior PDF of the damage condition properties, $\mathbf{z}\sim q_{\bm{\theta}^{*}}(\mathbf{z}|\mathbf{m},\mathbf{w})$. We assume that the unknown measurement error $\bm{\epsilon}$ is distributed according to a known PDF and that it is independent from the unknown damage condition $\mathbf{z}$ [32]. Thus, according to eq. (11), we can express the likelihood $p(\mathbf{m}|\mathbf{w},\mathbf{z})$ as:
$p(\mathbf{m}|\mathbf{w},\mathbf{z})=p\left(\mathbf{m}-\mathcal{F}_{\bm{\varphi}^{*}}([\mathbf{z},\mathbf{w}])\right)=p(\bm{\epsilon}).$ (12)
@@ 2.3.2 Loss function
Although $\mathbf{z}$ is unknown, its uncertainty can be modeled using a conditional probability distribution $p(\mathbf{z}|\mathbf{m},\mathbf{w})$. For any damage condition $\mathbf{z}$, the conditional PDF $p(\mathbf{z}|\mathbf{m},\mathbf{w})$ is the target posterior distribution that solves the inverse problem $\mathcal{I}$ under the Bayesian framework. Bayes’ theorem [82] implies that:
@@
Due to the intractability of the true posterior, we employ the inverse operator $\mathcal{E}_{\bm{\theta}}$ to estimate the parameters of an approximate PDF for each damage condition in $\mathbf{z}$, given the motion measurements $\mathbf{m}$ and operating loadings $\mathbf{w}$. Realizations from the approximate posterior $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ represent samples of the damage condition $\mathbf{z}$ that are likely produced by the unknown true posterior $p(\mathbf{z}|\mathbf{m},\mathbf{w})$. We remark that all the realizations $\{\mathbf{z}_{h}\}_{h=1}^{H}$ share the same operating conditions $\mathbf{w}$.
Our objective is to maximize the likelihood of the data while at the same time minimizing the discrepancy between the true posterior $p(\mathbf{z}|\mathbf{m},\mathbf{w})$ and the approximate posterior $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ obtained from $\mathcal{E}_{\bm{\theta}}$ To achieve this goal, we select the evidence lower bound (ELBO) as the loss function, a standard choice in VAE algorithms [83]. For any distributional model $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$, the evidence, $\log{p}(\cdot)$, is given by [25]:
$\begin{split}\log p(\mathbf{m},\mathbf{w})=\mathbb{E}_{q_{\bm{\theta}}}[\log p(\mathbf{m},\mathbf{w})]&=\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{p(\mathbf{m},\mathbf{w}|\mathbf{z})}{p(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]\\ &=\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{p(\mathbf{m},\mathbf{w}|\mathbf{z})q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})p(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]\\ &=\underbrace{\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{p(\mathbf{m},\mathbf{w}|\mathbf{z})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{ELBO}}}+\underbrace{\mathbb{E}_{q_{\bm{\theta}}}\left[\log\left[\frac{q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})}{p(\mathbf{z}|\mathbf{m},\mathbf{w})}\right]\right]}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}D_{\text{KL}}(q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})||p(\mathbf{z}|\mathbf{m},\mathbf{w}))}},\end{split}$ (14)
where $D_{\text{KL}}$ is the Kullback-Leibler divergence between two PDFs $p$ and $q$, and is defined as [25]:
$D_{\text{KL}}[p(x)||q(x)]=\int p(x)\ \text{log}\frac{p(x)}{q(x)}dx.$ (15)
Note that $D_{\text{KL}}[p||q]$ is zero if and only if $p=q$ [84]. Since $D_{\text{KL}}$ is non-negative by definition, the first term in eq. (14) constitutes a lower bound. Rearranging terms, we define the ELBO loss as:
@@
where we assume that the operating conditions $\mathbf{w}$ and the damage properties $\mathbf{z}$ are independent (i.e., $p(\mathbf{w},\mathbf{z})=p(\mathbf{w})\cdot p(\mathbf{z})$). Note that we can get rid of $p(\mathbf{w})$ in the second line of eq. (16) as it is independent of $\mathbf{z}$.
2.3.3 ELBO loss approximation
For a certain observation $\{\mathbf{m},\mathbf{w}\}$, we draw $H$ samples from the posterior and approximate the ELBO loss function via sample average approximation [85]:
@@ 2.3.3 ELBO loss approximation
The first term in eq. (17) accounts for the data misfit, which is the error between the true measurements and the reconstructions provided by $\mathcal{F}_{\bm{\varphi}^{*}}$. The second term refers to the prior, which we assume to follow a bounded uniform distribution $p(\mathbf{z})\sim\mathcal{U}[\mathbf{b}_{low},\mathbf{b}_{up}]$ with lower and upper bounds $\mathbf{b}_{low}$ and $\mathbf{b}_{up}$, respectively. The last term measures the probability that the $h$-th sample belongs to the estimated distribution $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$.
In this work, we assume the noise follows a Gaussian distribution, $p(\bm{\epsilon})=\mathcal{N}(0,\Gamma)$, where $\Gamma=\text{diag}(\beta\mathcal{F}_{\bm{\varphi}^{*}}([\mathbf{z},\mathbf{w}]))^{2}$ is a diagonal matrix containing the variances of the noise, and the parameter $\beta$ corresponds to the noise level. We can thus rewrite eq. (12) as:
By substituting the likelihood from eq. (18) into eq. (17), we finally express $\mathcal{L}_{\text{ELBO}}$ as:
@@
The second term in eq.19 can be neglected by directly constraining the Gaussian mixture density function to the desired interval.
@@
The next subsection explores different configurations for the approximate posterior $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ involved in eq.(20), including the necessary sampling strategy (see Figure 3).
@@ 2.4.1 Gaussian mixture
Let $q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$ be a Gaussian mixture (GM) of $K$ components such that:
@@
$\begin{gathered}q^{\text{GM}}_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w}):=\sum_{k=1}^{K}\alpha_{k}\mathcal{N}(\mathbf{z}|\bm{\mu}_{k},\Sigma_{k});\\ \mathcal{N}(\mathbf{z}|\bm{\mu}_{k},\Sigma_{k})=\frac{1}{(2\pi)^{D/2}|\Sigma_{k}|^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{z}-\bm{\mu}_{k})^{T}\Sigma_{k}^{-1}(\mathbf{z}-\bm{\mu}_{k})\right),\end{gathered}$ (21)
where $\alpha_{k}$ is the weight for the $k$-th Gaussian component, and $\mathcal{N}(\mathbf{z}|\bm{\mu}_{k},\Sigma_{k})$ is the corresponding multivariate Gaussian distribution with mean vector $\bm{\mu}_{k}$ and covariance matrix $\Sigma_{k}$. The Gaussian mixture posterior is currently defined over the entire space $\mathbb{R}^{D}$. However, the damage condition features $\mathbf{z}$ are bounded in a fixed domain $\mathcal{A}=[b_{\text{low}},b_{\text{up}}]$, where $b_{\text{low}}$ and $b_{\text{up}}$ stand for the lower and upper domain bounds, respectively. (maybe this is to be specified in the next section:)In this work, $\mathcal{A}$ corresponds to the $D$-dimensional unit hypercube $\mathcal{A}=[0,1]^{D}$. For compatibility with the Tensorflow libraries employed to implement this method, we truncate the Gaussian mixture by applying rejection sampling [86]. At each training iteration, we iteratively resample until the $H$ drawn samples belong to the domain, i.e., $\mathbf{z}^{h}\in\mathcal{A},\ h=1,...,H$. Algorithm 1 describes the rejection sampling step.
Algorithm 1 Truncated samples via rejection sampling
1:Unconstrained GM $q^{\text{GM}}_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$
2:Domain $\mathcal{A}$ with volume $\text{vol}(\mathcal{A})$
3:Number of samples $H$
4:A set of $H$ valid samples $\mathcal{S}\subset\mathcal{A}$
5:Draw initial batch $\mathcal{S}=\{\mathbf{z}^{1},\dots,\mathbf{z}^{H}\}$ from $q^{\text{GM}}_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$
6:Identify invalid indices $\mathcal{V}=\{j\mid\mathbf{z}^{j}\notin\mathcal{A}\}$
7:while $\mathcal{V}\neq\emptyset$ do
8:  Draw a new batch $\mathcal{S}_{\text{new}}=\{\mathbf{z}^{\prime 1},\dots,\mathbf{z}^{\prime H}\}$ from $q^{\text{GM}}_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$
9:  for each index $j\in\mathcal{V}$ do
10:   $\mathbf{z}^{j}\leftarrow\mathbf{z}^{\prime j}$
11:  end for
12:  Re-identify invalid indices $\mathcal{V}=\{j\mid\mathbf{z}^{j}\notin\mathcal{A}\}$
13:end while
14:return $\mathcal{S}$
To enforce the truncated samples to satisfy the PDF properties, we compute a normalization constant $C$ that accounts for the total probability mass of the unconstrained GM ($q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})$) that falls inside the target domain $\mathcal{A}$:
$C=\int_{\mathcal{A}}q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})d\mathbf{z}=\int_{b_{\text{low}}}^{b_{\text{up}}}q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})dz_{1}...dz_{D}$ (22)
The integral in eq. (22) is intractable. Therefore, we use Monte Carlo (MC) integration to approximate it:
$C=\int_{\mathcal{A}}q_{\bm{\theta}}(\mathbf{z}|\mathbf{m},\mathbf{w})d\mathbf{z}\approx\frac{\text{vol}(\mathcal{A})}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}q_{\bm{\theta}}(\mathbf{u}_{i}|\mathbf{m},\mathbf{w}),$ (23)
where $\text{vol}(\mathcal{A})$ corresponds to the volume of the finite domain $\mathcal{A}$.
@@
Previous works [52, 51] assumed diagonal covariance matrices. For any component in the mixture described in eq.(21), the covariance matrix $\bm{\Sigma}_{k}$ is formulated as $\bm{\Sigma}_{k}^{\text{diag}}=\text{diag}(\sigma_{1}^{2},...,\sigma_{Z}^{2})_{k}$, where $\sigma_{i}^{2}$ indicates the variance of the $i$th variable, and the subindex $k$ indicates the Gaussian component in the mixture (k=1,2,,Kk=1,2,...,K). For simplicity in notation, from now on we will omit the $k$ subindex to describe the covariance matrix specifications. The diagonal covariance matrix assumes a null interrelation between the involved variables $\mathbf{z}$, which enormously constrains the final shape of the mixture, hindering a fine approximation of the true solution [87].